OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4. 1 -1 -5 -25

OpenStudy (cggurumanjunath):

3x+3=-2x+2-4

OpenStudy (cggurumanjunath):

3x+3=-2x-2

OpenStudy (cggurumanjunath):

3x+2x=-2-3

OpenStudy (cggurumanjunath):

5x=-5

OpenStudy (cggurumanjunath):

x=(-5/5)